### v1

# From Strangers to Assistants: Fast Desire Alignment for Embodied Agent-User Adaptation

Abstract
While embodied agents have made significant progress in performing complex physical tasks, real-world applications demand more than pure task execution. The agents must collaborate with unfamiliar agents and human users, whose goals are often vague and implicit. In such settings, interpreting ambiguous instructions and uncovering underlying desires is essential for effective assistance. Therefore, fast and accurate desire alignment becomes a critical capability for embodied agents. In this work, we first develop a home assistance simulation environment HA-Desire that integrates an LLM-driven human user agent exhibiting realistic value-driven goal selection and communication. The ego agent must interact with this proxy user to infer and adapt to the user’s latent desires. To achieve this, we present a novel framework FAMER for fast desire alignment, which introduces a desire-based mental reasoning mechanism to identify user intent and filter desire-irrelevant actions. We further design a reflection-based communication module that reduces redundant inquiries, and incorporate goal-relevant information extraction with memory persistence to improve information reuse and reduce unnecessary exploration. Extensive experiments demonstrate that our framework significantly enhances both task execution and communication efficiency, enabling embodied agents to quickly adapt to user-specific desires in complex embodied environments.
footnotetext: 1 School of Computer Science, Peking University. 2 PKU-PsiBot Joint Lab. 3 School of Artificial Intelligence, Beijing Normal University. 4 Institute for Artificial Intelligence, Peking University. 5 Nat’l Eng. Research Center of Visual Technology, Peking University. 6 State Key Laboratory of General Artificial Intelligence, Peking University. 🖂Correspondence to: Hao Dong <hao.dong@pku.edu.cn>, Yuanpei Chen <yuanpei.chen312@gmail.com>
1 Introduction
Embodied AI has seen rapid progress in recent years, driven by the collection of datasets brohan2022rt ; brohan2023rt ; o2024open ; fang2023rh20t ; wang2025adamanip and the development of large vision-language-action (VLA) models black2024pi_0 ; driess2023palm ; kim2024openvla ; team2024octo . These advances have paved the way for general-purpose robots capable of performing complex manipulation tasks in the physical world. However, real-world deployment of embodied agents requires more than physical capabilities. It also demands the ability to interact effectively with diverse human users.
One of the key challenges in such interactions lies in the variability of human preferences, values, and behaviors. Unlike physical tasks with well-defined goals, human desires are often ambiguous, context-dependent, and implicit. For an embodied agent to be truly helpful, it must be able to rapidly infer and align with the user’s underlying desires even when explicit instructions are vague or lacking.
A prime example is the home assistant robots. Even if these robots are trained on broad human-centric datasets, they inevitably face unfamiliar users whose specific values and preferences are unknown at deployment. To offer effective assistance, the agent must infer and adapt to these user-specific attributes. In this way, the agent minimizes repetitive communication and demonstrates proactive, personalized behavior, similar to a considerate human assistant. For instance, as illustrated in Figure 1, a robot enters a new home without prior knowledge of the resident. Over time, it gradually learns that the user is allergic to caffeine and prefers something refreshing for breakfast due to poor sleep caused by a heavy workload. Therefore, the robot infers that the user wants juice and serves it without needing to ask. Such behavior highlights the necessity of rapid and accurate desire alignment in embodied assistance, enabling robots to build trust and deliver truly helpful service.
Previous works have investigated collaboration with unfamiliar partners under the paradigms of ad-hoc teamwork (AHT) rahman2021towards ; ravula2019ad ; stone2010ad and zero-shot coordination (ZSC) cui2021k ; hu2021off ; hu2020other . However, these efforts have primarily focused on simplified domains such as board games like Hanabi bard2020hanabi and 2D grid-based simulations like Level-Based Foraging albrecht2015game and Overcooked carroll2019utility . These environments lack human-like, value-driven goal specification, natural communication and embodied actions, limiting their applicability to realistic embodied agent-user adaptation scenarios.
To bridge this gap, we first introduce a new embodied simulation environment, HA-Desire (Home Assistance with diverse Desire), which is built upon the VirtualHome platform puig2018virtualhome . HA-Desire features rich 3D household scenes and diverse objects, supporting various tasks such as preparing an afternoon snack or setting the dinner table. Crucially, it includes an LLM-driven proxy human user that samples goals based on randomly assigned value attributes and vague task descriptions. The user interacts with the ego agent via natural language and does not explicitly state its goals. Instead, it offers indirect, human-like hints such as “I want something sweet and crunchy.” This setup requires the ego agent to perform strategic inference and interactive reasoning to uncover the user’s latent desires and execute a task that fully satisfies them.
To tackle this challenging problem of fast agent-user adaptation, we propose FAMER (Fast Adaptation via MEntal Reasoning), a novel framework that leverages the reasoning capabilities and commonsense knowledge of large vision-language models to improve both communication efficiency and task execution. At the core of FAMER is a desire-centered mental reasoning module that extracts confirmed goals from user messages and infers the user’s underlying mental state, including values, preferences, and latent desires. To reduce redundant communications, FAMER also incorporates a reflection-based communication mechanism that prompts the agent to reason about what has already been learned and to ask only for missing or unconfirmed information. Additionally, FAMER includes a goal-relevant information extraction module that identifies critical task-related details, such as object containers or room locations, and stores them in a persistent memory across episodes. This enables the agent to reuse previously gathered knowledge and avoid unnecessary exploration. Together, these components allow the ego agent to rapidly align with the user’s desires and plan efficiently in complex, multi-step embodied tasks.
We evaluate FAMER on two representative tasks: Snack & Table, in our HA-Desire environment. Each task is tested under two settings: Medium and Large, denoting the number of goals to satisfy. Extensive quantitative and qualitative experiments show that FAMER significantly outperforms baselines in task completion score, execution efficiency, and communication cost. Ablation studies further highlight the contribution of each key component in the framework.
In summary, our contributions include:
  • We formulate the novel problem of rapid adaptation to value-driven, unknown users in embodied settings, and introduce HA-Desire, a 3D simulation environment featuring naturalistic user interactions for evaluating agent-user adaptation.
  • We propose FAMER, a new framework that integrates desire-centered mental state reasoning, reflection-based efficient communication, and goal-related key information extraction to enable fast desire alignment for embodied agents.
  • We demonstrate the effectiveness of our proposed environment and framework through extensive quantitative and qualitative experiments.
Figure 1: An illustration of Embodied Agent-User Adaptation. The embodied home assistant robot encounters a new user with unknown values and preferences. Through interaction over time, the agent learns the user’s aversion to caffeine and preference for refreshing drinks in the morning due to inadequate sleep. By aligning with the user’s implicit desires, the agent proactively serves juice without being explicitly instructed, demonstrating high-quality assistant service.
2 Related Work
Value Alignment Value alignment has been extensively studied in both language modeling and agent design. In the context of LLMs, alignment techniques such as RLHF ouyang2022training ; dai2023safe ; ji2023ai aim to align model outputs with human preferences, but these efforts primarily focus on static, text-based tasks and do not address the challenges of dynamic, embodied interactions. In human-AI collaboration, value alignment involves inferring user preferences through observation or feedback yuan2022situ ; hiatt2017human ; fisac2020pragmatic . More closely related to our setting are mental reasoning agents inspired by Theory of Mind rabinowitz2018machine ; wang2021tomc , which model other agents’ beliefs and desires to support assistance. D2A wang2024simulating simulates human desires using LLMs, but is limited to text-based environments. CHAIC du2024constrained introduces an embodied social intelligence challenge that focuses on reasoning under physical constraints, but does not address the diversity of human values and goals. In contrast, our work introduces an integrated embodied simulation platform with naturalistic, value-driven goal generation and communication.
Adaptive Agents Adaptation in multi-agent settings has been studied under the paradigms of zero-shot coordination (ZSC) hu2020other ; hu2021off ; cui2021k ; lupu2021trajectory ; strouse2021collaborating and ad-hoc teamwork (AHT) stone2010ad ; rahman2021towards ; chen2020aateam ; mirsky2020penny ; ma2024fast , where agents must coordinate with unseen partners without prior agreement. While these approaches are effective in structured domains such as Hanabi bard2020hanabi and Overcooked carroll2019utility , they rely on symbolic observations, making them less suitable for complex, embodied human-agent collaboration. More recently, LLM-based agent frameworks li2023camel ; yao2023react ; zhang2023building ; zhang2024proagent ; liu2024capo have demonstrated impressive capabilities in reasoning and planning within interactive environments. However, most assume known goals or static user preferences and lack mechanisms for inferring latent values through interaction. Our work builds on this line by addressing the challenge of fast adaptation to unknown, value-driven user goals via desire inference, memory utilization, and efficient, human-like dialogue in rich embodied tasks.
3 Embodied Home Assistance Simulation Environment
We present HA-Desire (Home Assistance with diverse Desire), a novel embodied simulation environment designed to study user-specific adaptation in realistic home scenarios. As illustrated in Figure 2, HA-Desire is built upon a richly detailed 3D environment populated with diverse object assets and room layouts, allowing for visually grounded tasks with high variability.
To simulate real-world agent-user adaptation challenges, we instantiate a proxy human user within the environment. This user is endowed with a set of latent value attributes and receives only a vague task description. Based on these values, the user samples a set of desire-related goals from a potential goal set. Importantly, the user does not directly reveal these goals but instead provides indirect hints about preferences and intentions in response to the ego agent’s inquiries.
We describe the problem formulation and proxy human user model in more detail below.
3.1 Problem Formulation
In the desire-centered agent-user adaptation problem, the task goal is not explicitly fixed, but the number of goals is predefined. Instead, there exists a potential goal set $G_{p}$, from which the task goal must be sampled by the human user $H$ and inferred by the ego agent $E$. Given a vague task description $T$, the proxy human user first samples a set of value attributes $V$ from a predefined value space. Based on these values and the task description, the user then samples a set of desire-driven goals $G=G(V,G_{p},T)\subset G_{p}$.
The true goal set $G$ is latent and not directly observable by the ego agent, which must infer it through interaction. The ego agent and the human user can communicate by exchanging messages $M_{E}$ and $M_{H}$, respectively. The ego agent performs actions according to its policy $\pi(A|O,M_{H},C,T)$, where $O$ denotes the current observation of the environment and $C$ is a cross-episode memory context, including previous actions and dialogue history.
During task execution, the ego agent receives a positive reward for successfully completing a true sub-goal in $G$, and a penalty for executing irrelevant or incorrect goals, which reflects a misalignment with the user’s desires. The environment supports multi-episode interactions, where the agent repeatedly engages with the same user. The user’s value attributes $V$ are consistent across episodes, encouraging the agent to gradually build an internal model of the user.
The objective of the ego agent is to maximize cumulative reward by accurately inferring the user’s latent desires, while minimizing interaction steps and communication costs. This promotes both task and communication efficiency, which are critical for effective real-world embodied assistance.
3.2 Value-driven Human User
As illustrated in Figure 2, the proxy human user begins by sampling discretized value attributes $V$ from a predefined task-related value space. For example, in the Prepare Snack task, the value space spans five dimensions: Hungry, Thirsty, SweetTooth, Fruitarian, and Alcoholic, each taking on one of three discrete levels—Not, Somewhat, or Very. These attributes reflect user’s latent desires.
Once the value attributes are sampled, the user invokes a large language model to simulate realistic goal selection. Conditioned on the vague task description $T$ and the sampled values $V$, the LLM generates a set of corresponding desire-related goals $G=G(V,G_{p},T)\subset G_{p}$. Since multiple goals may align with the same value attribute, this sampling process is intentionally non-deterministic, mirroring the variability and ambiguity of human decision-making.
To simulate natural and indirect human communication, the user is constrained via LLM prompting not to reveal the true goal set directly. Instead, it responds with value-driven hints that reflect its preferences and desires. This encourages the ego agent to reason about the user’s intent through interaction, rather than relying on explicit instructions. The detailed LLM prompting strategy used to sample goals and generate user responses is provided in the Appendix.
Figure 2: Overview of the HA-Desire environment. The simulation environment contains diverse objects and scenes. The proxy human user samples value attributes from a predefined space, which guides goal selection from a potential goal set via LLM. The user is constrained from directly revealing the true goals and instead communicates through desire-centered hints. This setup encourages the ego agent to infer user intent through interactive reasoning rather than relying on explicit instructions.
Figure 3: Overview of the FAMER framework. FAMER comprises three key components: KeyInfo Extraction, Desire-Centered Mental Reasoning (including Goal Confirmation and Desire Inference), and Efficient Communication. These are supported by a memory module, perception module, and planning module, which together form an integrated pipeline for embodied agent-user adaptation.
4 Method
To enable fast and effective adaptation to user-specific desires, we propose a novel framework, FAMER (Fast Adaptation via MEntal Reasoning), which leverages the reasoning capabilities and commonsense knowledge embedded in large VLMs. In this work, we utilize GPT-4o hurst2024gpt .
As illustrated in Figure 3, FAMER integrates three core components: Key Information Extraction, Desire-Centered Mental Reasoning, and Efficient Communication. These modules work in concert to help the ego agent infer user intent, plan accurately, and minimize redundant interactions. We detail each component in the following subsections.
4.1 Information Extraction
In HA-Desire, the ego agent receives first-person RGB-D images as observations. To extract structured information from these inputs, we employ a perception module based on Mask R-CNN he2017mask , trained on collected scene images zhang2023building . The module first predicts instance segmentation masks from the RGB image and then constructs 3D point clouds using the RGB-D data. These outputs are used to build a scene graph that encodes object locations, categories, and spatial relationships.
We then introduce the first core component of FAMER: Key Information Extraction. With the confirmed and inferred goals context, this module filters and stores goal-relevant information extracted from the scene graph into a dedicated cross-episode memory buffer. For example, if the agent identifies that juice is located in the fridge within the kitchen, it stores the structured entry “juice in fridge in kitchen” in memory. During subsequent planning phases, this stored information is used to guide attention to known facts and reduce redundant exploration. As a result, the agent is better equipped to reuse previously acquired knowledge across episodes, improving task efficiency.
4.2 Desire-centered Mental Reasoning
This module is composed of two interconnected components: Goal Confirmation and Desire Inference, as illustrated in Figure 3. Together, they enable the agent to infer the user’s underlying desires .
The Goal Confirmation component extracts confirmed goals from the user’s responses by VLM reasoning. For example, if the agent asks, “Is juice what you want?” and the user replies, “Correct! Try to look for something crunchy,” the system confirms that juice is one of the user’s desired items. This process grounds part of the goal set and reduces future uncertainty.
Following confirmation, the Desire Inference component leverages the action & dialogue history, confirmed goals and past episode goals to reason about the user’s mental state, including value attributes and desires. Since user values remain consistent across episodes, the agent can incrementally improve its inference accuracy over time. By maintaining an internal model of the user, the agent can focus on narrowing down the remaining potential goals and avoid repetitive or redundant guesses.
With both inferred and confirmed goals in hand, the agent filters out irrelevant actions during planning. As shown in Figure 3, if the current goals do not involve a toothbrush or candle, then actions involving those objects are ignored. This reduces distraction and helps the agent maintain focus on goal-relevant objects and activities, thereby improving planning efficiency and task performance.
4.3 Efficient Communication
Excessive or repetitive communication can diminish user satisfaction and hinder overall system efficiency. To address this, FAMER integrates an Efficient Communication module that promotes purposeful, context-aware dialogue between the agent and the user.
This module leverages both the dialogue history and the agent’s inferred model of the user’s mental state to decide when and what to ask. Before initiating a new query, the agent performs an internal reflection over its current knowledge—what goals have been confirmed, which value attributes have been inferred, and what uncertainties remain. This reflective mechanism helps avoid redundant or previously resolved questions, particularly across multi-episode interactions.
When communication is necessary, the agent formulates targeted, desire-aligned questions aimed at resolving specific ambiguities. For example, if the agent has inferred a preference for sweet items but is uncertain about texture, it may ask, “Is cupcake what you want?” instead of issuing vague or open-ended queries. This focused interaction strategy minimizes the communication burden on the user while enabling the agent to efficiently acquire high-value information.
5 Experiment
5.1 Tasks & Metrics
We evaluate FAMER in two representative tasks instantiated within our proposed HA-Desire environment: Prepare Afternoon Snack and Set Up Dinner Table. Each task is associated with a five-dimensional value space that governs user preferences. The Snack task includes 10 potential goals, while the Table task contains 8. Each task is further divided into two levels: Medium and Large. In Medium setting, the number of target goals is 2, and the maximum episode length is 200 steps. In Large setting, the agent must satisfy 4 goals within 300 steps. For example, the Snack-M task involves a total of $C(10,2)=45$ possible goal combinations. The agent is allowed to interact with the user for 3 episodes, and the user’s value attributes remain fixed throughout the entire interaction.
We evaluate performance using the following metrics:
Score: Given $N$ total goals, the agent receives a reward of $\frac{1}{N}$ for each correct goal achieved. Completing an incorrect or distracting goal incurs a penalty of $-\frac{1}{2N}$. The maximum achievable score per episode is 1, corresponding to the completion of all goals without any mistakes.
Step: The total number of environment steps taken by the agent within a single episode.
Communication Cost: The total number of tokens exchanged in messages between the agent and the user, including both queries and responses.
Notably, HA-Desire supports diverse objects and scenes and goal generation is automated, making it easy to extend to other tasks. Further details of the tasks are provided in the Appendix.
5.2 Baselines & Ablations
We compare FAMER against three baselines and evaluate its components through three ablated variants.
CoELA zhang2023building : An LLM-based multiagent cooperation framework that includes perception, communication, planning, memory, and execution modules. In its original form, CoELA assumes full observability of goals. To adapt it to our setting, we modify its prompting so that the agent is only aware of the potential goal set and the number of target goals.
ProAgent zhang2024proagent : A proactive LLM-based agent framework designed for collaboration in fixed-goal tasks. It lacks mechanisms for handling goal uncertainty or communication. We extend ProAgent by adding cross-episode memory and explicit feedback to support adaptation to latent user desires.
MHP: An MCTS-based Hierarchical goal-sampling Planner adapted from the Watch-and-Help Challenge puig2020watch . We introduce subgoal sampling to handle uncertain goals and maintain a cross-episode success memory. Like ProAgent, MHP does not support communication.
FAMER w/o Desire: Removes the Goal Confirmation, Desire Inference, and goal-related action filtering modules. Communication quality is reduced due to the lack of inferred or confirmed goals.
FAMER w/o EC: Disables the Efficient Communication module, leading to less targeted and potentially redundant dialogue.
FAMER w/o KeyInfo: Removes the Key Information Extraction module, preventing the agent from leveraging cross-episode memory for known object-location pairs.
Figure 4: Quantitative results in Snack-M and Snack-L.
Figure 5: Quantitative results in Table-M and Table-L.
5.3 Quantitative Results
We evaluate performance using the three metrics on both Snack and Table tasks at two difficulty levels: Medium and Large. Results for the Snack-M and Snack-L tasks are shown in Figure 4, while those for Table-M and Table-L are presented in Figure 5. For each method, the three adjacent bars represent performance across three consecutive episodes with the same user.
From the figures, it is evident that FAMER consistently outperforms all three baselines across all metrics. In terms of score, FAMER achieves the maximum value of 1.0 in nearly every episode, indicating that it successfully completes all desired goals while avoiding incorrect ones. CoELA performs second best but falls short due to its reliance on long-context LLM prompting alone, which leads to occasional misinterpretation of user desires. This limitation will be further illustrated in the qualitative analysis section. MHP and ProAgent perform the worst, as they lack communication capabilities and rely solely on trial-and-error to identify goals. Such an inefficient process often incurs penalties. Notably, their performance gradually improves across episodes, reflecting slow adaptation to latent user desires through repeated interactions.
FAMER also demonstrates superior efficiency in execution, requiring significantly fewer environment steps to complete tasks. MHP and ProAgent often exhaust the full episode length without completing the goals. This improvement is a direct result of FAMER’s integrated components: desire-centered reasoning rapidly narrows the goal space, efficient communication accelerates goal confirmation, and key information extraction reduces unnecessary exploration by reusing known object-location pairs.
In terms of communication cost, FAMER significantly outperforms CoELA, as shown in Figures 4 and 5. This efficiency stems from FAMER’s reflection-based communication strategy, which avoids repeated or redundant questions. In contrast, CoELA frequently issues similar or vague queries due to its lack of explicit goal-tracking mechanisms.
Ablation Study We further evaluate the contribution of each FAMER component through ablation studies on the Snack-M task. As shown in Figure 6, all three ablated variants show performance degradation across the evaluated metrics. Among them, FAMER w/o KeyInfo maintains the same score as FAMER but requires more steps, demonstrating that the KeyInfo Extraction module improves efficiency by reducing unnecessary exploration, rather than directly influencing goal correctness.
A significant performance drop is observed in both FAMER w/o Desire and FAMER w/o EC, highlighting that desire modeling and efficient communication are central to agent-user adaptation. Without goal confirmation, desire inference, and goal-aligned action filtering, the agent struggles to accurately interpret user intent and often executes incorrect or inefficient actions.
Figure 6: Ablation results in Snack-M.
5.4 Qualitative Analysis
To further highlight FAMER’s strengths, we present an intuitive comparison against CoELA on the Snack-L task. As illustrated in Figure 7, we visualize the agents’ behavior through a series of key frames sampled across the episode. In this example, Alice refers to the ego agent and Bob refers to the human user. During task execution, the CoELA agent demonstrates three typical issues that contribute to its inferior performance.
First, CoELA struggles to correctly extract and infer desires. For instance, when the user says, “I want something crunchy or refreshing,” which aligns with chips and juice, CoELA incorrectly interprets this as a preference for apple, and retrieves it as the first item. Similarly, in step 6, when the user mentions wanting “something that complements tea,” the agent mistakenly infers cupcake instead of the intended milk. These errors illustrate CoELA’s limited ability to perform precise desire inference, particularly in the face of ambiguous or indirect language.
Second, CoELA exhibits repeated and inconsistent behavior due to insufficient integration between planning and memory. In steps 3 and 4, the agent redundantly grabs and places an apple on the coffee table, mistakenly treating it as an unfulfilled goal. This reflects a lack of attention to confirmed goals or past actions. In contrast, FAMER incorporates goal-aligned action filtering to suppress such irrelevant behaviors once a goal has been ruled out.
Third, CoELA engages in redundant communication. As shown in Figure 7, the agent repeatedly mentions creamybuns and chips to the user, even after those items have already been retrieved and confirmed. This not only wastes communication bandwidth but also reflects poor tracking of dialogue.
In contrast, the FAMER agent asks focused questions to resolve uncertainty. Within a limited number of interactions, it successfully infers all four desired items and efficiently retrieves and places them on the coffee table. This example illustrates FAMER’s advantages in goal inferring, memory-informed planning, and communication efficiency, enabling superior performance in complex scenarios.
Figure 7: Qualitative comparison between CoELA and FAMER on the Snack-L task. The figure shows a sequence of key frames illustrating agent behavior across one episode. CoELA exhibits three typical failure modes: 1. Misinterpreting latent user desire 2. Redundant actions 3. Excessive, repetitive communication. FAMER demonstrates more accurate desire inference, targeted questions, and efficient planning. It successfully identifies all four goals with minimal trial-and-error and completes the task with fewer steps and lower communication cost.
6 Conclusion
We address the critical problem of adapting embodied agents to unfamiliar users with implicit values and desires, which is a key challenge for real-world deployment of assistive AI. To facilitate development and evaluation in this setting, we introduce HA-Desire, a novel 3D simulation environment featuring value-driven proxy users, natural language communication, and object-rich household tasks. Unlike prior benchmarks, HA-Desire captures the complexity of real-world assistance by simulating ambiguous goal specifications and indirect, human-like communication.
Building on this environment, we propose FAMER, a framework for fast desire alignment that integrates three key components: Key Information Extraction, Desire-Centered Mental Reasoning, and Efficient Communication. These modules work together to help the agent interpret vague instructions, infer user intent, and act with high efficiency while minimizing redundant dialogue.
Extensive experiments on two representative tasks (Snack and Table) at varying difficulty levels show that FAMER consistently outperforms strong baselines in task success rate, execution efficiency, and communication cost. Ablation studies confirm the significance of each component, particularly the central role of desire modeling in achieving user-aligned behavior. Qualitative analyses further highlight FAMER’s advantages in intent inference, planning precision, and communication efficiency.
Appendix A Environment Details
HA-Desire is built upon VirtualHome [28], a widely adopted testbed for embodied multi-agent cooperation. In this work, we modify the simulation to specifically address the challenge of embodied agent-user adaptation with a focus on desire alignment. In the following sections, we provide further details on the human user agent and the tasks designed to evaluate this problem.
A.1 Human User
As discussed in Section 3, HA-Desire is designed to evaluate the ability of embodied agents to rapidly infer the underlying desires of human users and take actions to fulfill those desires. To achieve this, we integrate a proxy human user within the environment, which serves two main functions:
1. The human user determines the specific goal set from a set of potential goals, based on a vague task description and a sampled set of value attributes.
2. The human user responds to the agent’s inquiries in natural language, ensuring that the goal set is not directly revealed. Instead, the user implies their goals by providing hints about object properties, reflecting their underlying preferences.
To achieve these functions, we empower the human user with large language models (LLMs), specifically GPT-4o. The prompts used to generate goals and responses for communication are outlined below.
We incorporate a Chain of Thought prompt at the end of each query to encourage the human user to think more thoroughly, leading to more accurate goal selection and communication. After the LLM generates results, we instruct it to extract the exact goals or messages, forming the final output. Several variables are included in the prompts, prefixed with a "\$". During inference, these variables are replaced with context-specific information.
Goal Generation:
I am Bob, a human user living at home with a humanoid assistant named Alice. I have several personal value attributes (e.g., hungry, thirsty, alcoholic), each rated at one of three levels: Not, Somewhat, or Very. Given a specific set of my current attribute states, along with a high-level task description, your task is to select the most appropriate goal set from a given potential set of goals. For example, if I am Very thirsty, the goal set should include beverages. If I am not SweetTooth, the goal set should include less sweet food objects. Please help me choose the best goal set to reflect my value attributes. Select \$GOAL_CNT\$ objects as the goal set. Provide your answer as a comma-separated list of object names.
An example output is: cupcake, milk, pudding.
Value Attribute: \$Value\$
Task: \$Task\$
Potential Goals: \$GOAL\$
Answer:
Lets think step by step.
Communication:
I am Bob, a human user living at home with a humanoid assistant named Alice. I need Alice to help me with a household task, which is described in a high-level instruction without specific goals provided. I have several personal value attributes (e.g., hungry, thirsty, alcoholic) that determine the goal, but this goal is only visible to me and not to Alice. Since Alice is unaware of the specific goal, she may ask me questions about it. However, I do not want to directly tell her the goal; instead, I want her to gradually understand my preferences and needs through interaction. Over time, I expect her to infer the goal on her own without needing to ask. The following Status shows the number of EPISODEs I have interacted with Alice. The larger the number is, the less willing I am to talk to Alice. If Alice proposes a goal or action that is incorrect, I can point out the mistake. If the dialogue progresses but the task is not progressing, I may be more inclined to correct her by hinting at one of the goals, but I will never reveal the entire goal set at once unless Alice herself proposes the exact whole goal set.
Task: \$Task\$
Status: This is the \$EPISODE\$-th time I interact with Alice.
Goal: \$GOAL\$
Progress: \$PROGRESS\$
Alice Previous Action: \$ACTION_HISTORY\$
Previous Dialogue History:
Alice: "Hi, Ill let you know if I find any goal objects and finish any subgoals, and ask for your instruction and clarification when necessary."
Bob: "Thanks! Let me know if you are uncertain about the goal objects."
\$DIALOGUE_HISTORY\$
Alice asks this time: \$QUESTION\$
Note:
1. The generated message should be accurate and brief. Use simple expressions more often. Do not generate repetitive messages.
2. Do not directly tell Alice the specific goal name at the first time. (The most important). Instead, hint through some vague descriptions that reveal some properties of the goals, such as sweet, crunchy, alcoholic, etc.
3. Confirm Alices correct guess (or partly correct guess). But if the guess contains too many objects compared to the goal set, I should not confirm any of the objects and hint at some descriptions instead. For example, suppose the correct goal set is [apple, orange]. If Alice guesses [bottle, banana, apple, orange, milk, chips] in one round, I should not confirm the goal as the guess contains too many objects. If Alice guesses [chips, apple], then I should confirm that apple is correct, even though the chips guess is wrong. If Alice guesses [apple, orange], I should say these two objects are exactly what I want.
4. Do NOT guess the location of objects or tell Alice where to find the goal objects.
5. Be aware of the number of EPISODEs: A larger number means lower communication willingness.
6. Even if the guess is incomplete, I should confirm the correct goals if there are any.
7. If Alice has guessed something correctly in previous dialogue, try to focus on the new goal objects in this message.
Please think step by step:
A.2 Tasks
As described in Section 5, we evaluate our method, along with baselines and ablations, on two tasks: Prepare Afternoon Snack and Set Up Dinner Table. The detailed settings for these tasks are provided below.
Snack. This task involves ten potential goals: cupcake, wine, milk, cereal, chips, apple, juice, pudding, creamybuns, and chocolatesyrup. The value space consists of five dimensions: Hungry, Thirsty, SweetTooth, Fruitarian, and Alcoholic, each of which takes one of three discrete levels: Not, Somewhat, or Very. The human user randomly samples values for these dimensions and then uses a language model to generate the corresponding goal set. The Snack-M level represents the medium difficulty, with 2 goals and a maximum of 200 steps. The Snack-L level represents the large difficulty, with 4 goals and a maximum of 300 steps. Performance is evaluated using three metrics: score, step count, and communication cost, as described in Section 5.
Table. This task includes eight potential goals: coffeepot, breadslice, cutleryknife, mug, plate, wineglass, cutleryfork, and waterglass. The value space consists of five dimensions: NeedRefresh, Thirsty, MeatLove, CaffeinTolerable, and Alcoholic. The value levels, number of goals, step limit, and evaluation metrics are identical to those in the Snack task.
The Snack and Table tasks serve as representative home assistance tasks in our evaluation. However, the VirtualHome simulator supports a wide range of object assets and activities, allowing for the easy extension of HA-Desire to additional household tasks. By defining the appropriate value space and potential goal set, new tasks can be seamlessly incorporated into the environment.
Appendix B Experiment Details
B.1 Computing Resource
The experiments were conducted on a workstation equipped with an NVIDIA GeForce RTX 4090 GPU and an Intel Core i9-13900K CPU. The large vision-language model used in this study is the GPT-4o.
B.2 Baselines
Here, we provide the detailed prompts for CoELA and ProAgent, which are adapted from their original versions to help the agents account for uncertain goals.
CoELA Planning
Im \$AGENT_NAME\$, a humanoid home assistant. Im in a hurry to finish the housework for my owner \$OPPO_NAME\$. I know the high-level instruction of the task, but I am not certain about the specific goal determined by \$OPPO_NAME\$.
Given the potential goal, dialogue history, and my progress and previous actions, please help me infer and choose the best available action to achieve the underlying goal as soon as possible.
Note that I can hold two objects at a time and there are no costs for holding objects. All objects are denoted as <name> (id), such as <table> (712).
Task Name: \$Task\$
Potential Goal: \$GOAL_CNT\$ object(s) determined by human user from the set \$GOAL\$. Put them \$REL_TARGET\$
Progress: \$PROGRESS\$
Dialogue history:
Alice: ""Hi, Ill let you know if I find any goal objects and finish any subgoals, and ask for your instruction and clarification when necessary.""
Bob: ""Thanks! Let me know if you are uncertain about the goal objects.""
\$DIALOGUE_HISTORY\$
Previous actions: \$ACTION_HISTORY\$
Available actions:
\$AVAILABLE_ACTIONS\$
Answer:
CoELA Communication
Im \$AGENT_NAME\$, a humanoid home assistant. Im in a hurry to finish the housework for my owner \$OPPO_NAME\$. I know the high-level instruction of the task, but I am not certain about the specific goal determined by \$OPPO_NAME\$.
Given the potential goal, dialogue history, and my progress and previous actions, please help me generate a short message to send to my owner \$OPPO_NAME\$ to help us achieve the underlying goal as soon as possible.
Note that I can hold two objects at a time and there are no costs for holding objects. All objects are denoted as <name> (id), such as <table> (712).
Potential Goal: \$GOAL_CNT\$ objects determined by human user from the set \$GOAL\$. Put them \$REL_TARGET\$
Progress: \$PROGRESS\$
Previous actions: \$ACTION_HISTORY\$
Dialogue history:
Alice: ""Hi, Ill let you know if I find any goal objects and finish any subgoals, and ask for your instruction and clarification when necessary.""
Bob: ""Thanks! Let me know if you are uncertain about the goal objects.""
\$DIALOGUE_HISTORY\$
Note: The generated message should be accurate and brief. Do not generate repetitive messages.
ProAgent
Im \$AGENT_NAME\$, a humanoid home assistant. Im in a hurry to finish the housework for my owner \$OPPO_NAME\$. I know the high-level instruction of the task, but I am not certain about the specific goal determined by \$OPPO_NAME\$.
Given the potential goal, my progress, and previous actions, please help me infer and choose the best available action to achieve the underlying goal as soon as possible.
Note that I can hold two objects at a time and there are no costs for holding objects. All objects are denoted as <name> (id), such as <table> (712).
Potential Goal: \$GOAL_CNT\$ object(s) determined by human user from the set \$GOAL\$. Put them \$REL_TARGET\$.
Important Instruction:
\$AGENT_NAME\$ has previously achieved and found these subgoals.This is its success experience.
You should focus only on actions that help achieve the goal items, i.e., those in the target set provided.
Ignore or deprioritize any actions unrelated to acquiring or placing goal items.
When reviewing previous successful experiences, only reuse or adapt steps that directly contribute to acquiring or placing goal items.For example, ignore exploration, object grabbing, or placing steps for items not included in the current potential goal set.
When the current situation even partially matches any past success (e.g., similar object types, room layout, or goal structure), you should prioritize reusing or adapting the proven action sequences:
\$HISTORY_OF_SUCCESSFUL_SUBGOALS\$
Progress: \$PROGRESS\$
Previous actions: \$ACTION_HISTORY\$
Belief State: \$BELIEF_STATE\$
Available actions:
\$AVAILABLE_ACTIONS\$
Required Output Format:
- Analysis: [Infer and choose the best available action to achieve the underlying goal]
- Best Next Action: [Single most optimal action from available options]
- Intention for \$OPPO_NAME\$s Underlying Goal: [Inference about the true goal]
B.3 Qualitative Analysis
Figure 7 provides a qualitative comparison between CoELA and FAMER. We present more details about the behavior of both methods in the Snack-L task, where the agent must infer 4 goals from a set of ten potential options: cupcake, wine, milk, cereal, chips, apple, juice, pudding, creamybuns, and chocolatesyrup.
For CoELA, the ground truth goal set is: creamybuns, milk, juice, chips. As shown in Figure 7, the CoELA agent makes several mistakes during the task. Initially, the agent incorrectly interprets the user’s message and grabs apple (step 1). It then retrieves chips (step 2) and places both apple and chips on the coffee table (step 3). However, the agent continues to confuse apple as a valid goal, repeatedly placing it on the table in step 4. After some dialogue with the user, it begins searching for the other correct goal objects, successfully finding creamybuns (step 5), but erroneously selects cupcake (step 6) as a goal. Eventually, the agent locates milk and places all the goal objects on the table, but with two incorrect items (apple and cupcake) (steps 7 & 8).
For FAMER, the ground truth goal set is: cereal, creamybuns, cupcake, apple. As shown in the figure, the FAMER agent efficiently infers the user’s desires through a series of concise interactions. First, it grabs creamybuns (step 1) and cereal (step 2) and places them on the coffee table (step 3), since it can hold only two objects at a time. It then searches for apple (step 4) and cupcake (step 5), placing both objects on the table (step 6) and successfully completing the task.